### v1

# Non-Markovianity of subsystem dynamics in isolated quantum many-body systems

Abstract
It is believed that an isolated quantum many-body system far away from equilibrium should try to attain equilibrium via a mechanism whereby any given subsystem acts as an open quantum system that is coupled to an environment which is the complementary part of the full system and undergoing a complicated equilibration process such that all the subsystems in the long-time limit attain equilibrium states compatible with the global equilibrium state. This picture begs the question whether the dynamics of any given subsystem is Markovian (monotonic loss of information and memory) or non-Markovian. In this work, by numerically probing the dynamical behaviour of the distance between temporally-separated subsystem states, we reveal the telltale signatures and other associated features of quantum (non-)Markovianity of the dynamics of small subsystems of an isolated quantum spin system in one dimension (the Ising spin chain) brought far from equilibrium by quantum quenches. Additionally, remarkably systematic behaviour is seen in the dynamics of the eigenvalues of the reduced density matrices of the subsystems. These features strongly depend on the direction of quenching in the parameter space, with paramagnetic-to-ferromagnetic quenches offering considerably stronger signatures of subsystem non-Markovianity, for which we offer heuristic arguments.
Introduction— It is expected that a closed quantum many-body system when brought far out of equilibrium by quenching one of its parameters will after a sufficiently long-time equilibriate to a state as dictated by the post-quench parameters. The actual process of equilibration can be rather complex because during this process the system will explore the set of accessible states (decided by the extant conserved quantities) in the Hilbert space during its Hamiltonian dynamics, and the complexity of the trajectories in the quantum state space depends crucially on the initial state and its relation with the Hamiltonian driving its evolution. For generic (i.e., non-integrable) systems with only the total energy being a conserved quantity, the long-time physics should then be described by usual statistical mechanics in the microcanonical picture, which essentially encapsulates the notion of ergodicity. One then says that the system has thermalized, i.e., the long-time (thermal) state is the Gibbs density matrix as described by the Gibbs ensemble, and the eigenstate thermalization hypothesis (ETH) describes well the behaviour of observables in this regime [1, 2, 3, 4]. Exceptions to this paradigm within the class of non-integrable quantum many-body systems exist of course [5, 6, 7, 8]. For integrable quantum many-body systems, i.e., those with an extensive number of conserved quantities, the long-time physics is instead described by a so-called generalized Gibbs ensemble picture wherein the equilibrated state is a generalized Gibbs density matrix [9, 10], and one says that these systems do not thermalize in the usual sense (but possibly in a generalized sense [11]) and constitute a strong violation of ergodicity and the ETH.
However, the nature and characteristics of the Hamiltonian dynamics in the far-from-equilibrium regime (early and intermediate times), where organizing principles or structures are scarce, is much less understood. In particular, it seems prudent to have a broader classification of general characteristics of the dynamics that goes beyond the existing classifications such as fast-, slow- or non-thermalizing/equilibrating as measured by e.g. the growth of entanglement entropies or the behaviour of local observables with time. One natural way to think about this dynamical process is by considering any subsystem of the full system as an open quantum system interacting with its environment, i.e., its complementary part of the full system, and it is natural to presuppose that it is in fact the nature of the interactions between the various subsystems acting as environments for each other that decides the various facets of the non-equilibrium relaxation dynamics.
With this line of thinking, and employing the notion of distances or divergences between quantum states of which many measures exist in (quantum) information theory [12, 13, 14], our work investigates the following basic question : how does the distance between temporally separated quantum states of a given subsystem evolve in time and what underlying characteristics of the non-equilibrium dynamics are thereby revealed. Independently, the viewpoint of regarding a given subsystem as an open quantum system necessitates asking whether its dynamics is Markovian (memory-less, and monotonic loss of information) or not [15, 16, 17, 18]. Both these questions are intimately related to each other, and we believe this classification of subsystem dynamics as Markovian or non-Markovian is a concrete addition to the existing list of characteristics of the non-equilibrium dynamics of closed quantum many-body systems.
Quantum states and channels— Herein we introduce our notions and notations. Quantum states live in the Hilbert space $\mathcal{H}$, which for quantum spin systems on $N$ sites assumes a tensor product structure $(\mathbb{C}^{d})^{\otimes N}$ over the $N$ local Hilbert spaces each of dimension $d$ (e.g. $d=2$ for $S=1/2$ spins). Often though by quantum states we shall refer to the density matrix (operator) $\rho$. It is Hermitian and positive semi-definite with unit trace, which means that its eigenvalues $\{p_{i}\}$ can be regarded as providing a bonafide probability distribution. The set of density matrices is convex and will be denoted by $\mathcal{D}(\mathcal{H})$.
A trace-preserving quantum operation $\Lambda$ is called positive if $\Lambda\!:\!\mathcal{D}(\mathcal{H})\!\rightarrow\!\mathcal{D}(\mathcal{H})$ (this holds more generally over the space of linear and bounded Hermitian operators, of which $\mathcal{D}(\mathcal{H})$ is a particular subset). A more powerful and non-trivial notion is complete-positivity, which means that $\mathbb{I}\otimes\Lambda$ is positive, where the identity operation $\mathbb{I}$ is understood to act on those parts of the system that are complementary or ancillary to the support of the action of $\Lambda$. Completely-positive trace-preserving (CPTP) maps are also known as quantum channels. Examples include unitaries, measurements and partial trace. Physically (meaning, experimentally) realizable operations have to be CPTP.
Distances between quantum states— We choose the trace distance as our primary measure of distance between quantum states in this article (results are qualitatively unchanged if other measures are chosen). Given two density matrices $\rho$ and $\sigma$, the trace distance (TD) between them is defined as $T_{d}(\rho,\sigma)\!=\!\frac{1}{2}\sum|\lambda_{i}|$, where $\{\lambda_{i}\}$ is the set of eigenvalues of $(\rho-\sigma)$, with the index $i\!=\!1,2,..,\operatorname{rank}(\rho-\sigma)$ The trace distance takes values $\in[0,1]$. Let $\{p^{\downarrow}\}$ and $\{q^{\downarrow}\}$ be the vector of eigenvalues of $\rho$ and $\sigma$ respectively, arranged in a descending order. The counterpart of trace distance for probability distributions, known as the total variation distance (TVD), is defined as $V_{d}(p^{\downarrow},q^{\downarrow})\!=\!\frac{1}{2}\sum_{i}|p^{\downarrow}_{i}-q^{\downarrow}_{i}|$, and it also takes values $\in[0,1]$.
Divisibility, information backflow and non-Markovianity— A fundamental and powerful property of any distance measures is that they are non-increasing (or "contractive") under CPTP maps [12, 13, 14, 19]. That is, given a CPTP operation $\Lambda:\mathcal{D(\mathcal{H})}\!\rightarrow\!\mathcal{D(\mathcal{H})}$, one has $T_{d}(\Lambda(\rho),\Lambda(\sigma))\leq T_{d}(\rho,\sigma)$ for the trace distance. In other words, any (physically realizable) quantum operation can not increase the (information about the) distance between (and hence, a quantifier of "distinguishability") of two quantum states.
Consider now a dynamically evolving quantum system. Let $\Lambda_{t,0}$ denote the family of CPTP maps that evolves the subsystem from initial time $t=0$ to time $t$ for all $t>0$, these will be referred to as dynamical maps. This temporal evolution process is said to be (CP-)divisible if one has $\Lambda_{t,0}=\Lambda_{t,s}\Lambda_{s,0}$ for any $s\in(0,t)$, where $\Lambda_{t,s}$ is also a CPTP map evolving the system from time $s$ to $t$, otherwise if $\Lambda_{t,s}$ is not CPTP, the evolution is termed (CP-)indivisible (note that it is not to be taken for granted nor is it obvious or generic that the intermediate map $\Lambda_{t,s}$ has to be CPTP). It then follows that in a divisible process, the aforementioned monotonicities are respected at every step of the process, and so ultimately one has $T_{d}(\Lambda_{t,0}(\rho),\Lambda_{t,0}(\sigma))\leq T_{d}(\rho,\sigma)$ for any $t>0$. The distance between two quantum states continues to decrease all along the temporal dynamics, and such a dynamical process is termed Markovian, and any violation of this condition can be taken as a signature of non-Markovianity [20, 15, 16]. Thus, an CP-indivisible process is non-Markovian though the converse does not necessarily hold true [21]. An alternative approach is to view the upholding of these monotonicities as monotonic loss of information ("memory") about the distinguishable features of a pair of quantum systems coupled separately to the same environment (which is stable by itself and is ideally not affected by its interactions with the non-equilibrium quantum systems to which it is coupled, which is the essence of the Born-Markov approach to open quantum systems [22, 23]), whereby the environment acts to drive the systems to the final equilibrium states and thereby decreasing any distinguishable features (absorbing the information about such distinguishabilities) between them over time; such a process is Markovian by this token [16]. Consequently, any violation of these monotonicities signal backflows of information from the environment to the systems (i.e., the environment somehow retained the dynamical memory of the past states of the systems; the system-environment coupling is beyond the Born-Markov regime and the environment is in fact itself affected by the systems it is coupled to and unable to quickly shed information about them) and this is a defining signature of non-Markovianity [24, 25]. These two viewpoints are not necessarily equivalent to each other in general nor is it obvious that they should be, however for invertible CPTP maps [[Invertibilityisnottobeconfusedwithreversibility, andinvertiblemapsarenotgenerallyCPTPeither.OnlyforaunitaryCPTPmapisitsinversealsoCPTPandconsequentlyinvertibilityisthesameasreversibility, seefore.g.][.]Chruscinski2022a] (that is, $\Lambda_{a,b}^{-1}$ exists, meaning none of the eigenvalues of $\Lambda_{a,b}$ is zero; such is the case for us as explicitly checked numerically) these two viewpoints are understood to be equivalent [27, 28, 29, 18]. To be sure, our numerical results are really only measuring the backflows of information, and not actually establishing if all the intermediate $\Lambda_{t,s}$ are CPTP or not and thereby establishing (in)divisibility from first principles, which is a separate issue in itself and is being addressed in a parallel work. It is only thanks to these maps being invertible that we talk of both these notions of non-Markovianity interchangeably in this work.
Consider then a subsystem (not necessarily contiguous) of length $l$ spins, represented by a density matrix $\rho_{t}^{(\ell)}$ at time $t$, of a non-equilibrium quantum many-body system that is evolving under its own Hamiltonian ($t=0$ denotes the starting time). We shall be dealing exclusively with translationally invariant systems in this work, thus it does not matter where exactly this $-\ell$ subsystem is located within the full system as long as it is located in the bulk away from the boundaries (in case of open boundary conditions). We probe the following basic questions in this work : how the distance between two temporally-separated states of a subsystem behaves as the system evolves in time, and what characteristics of the non-equilibrium dynamics is revealed by this behaviour, and if there are any invariants or qualitative changes under varying temporal separation of considered subsystem states ? As such, our main quantities of interest are $T_{d}(\rho_{t+\delta}^{\ell},\rho_{t}^{\ell})$, where $\delta$ is the temporal separation. Let $\Lambda_{[t]}$ be the dynamical CPTP map that evolves the subsystem by time $t$, i.e., $\rho_{t+a}^{\ell}\!=\!\Lambda_{[t]}[\rho_{a}^{\ell}]$, then by previous discussion, if the dynamics is divisible, i.e., $\Lambda_{[t]}\!=\!(\Lambda_{[1]})^{t}$ (more generally, in terms of the previous discussion, $\Lambda_{[t]}\!=\!\Lambda_{t+a,a}$ for any reference time stamp $a$), then a non-increasing behaviour of $T_{d}(\rho_{t+\delta}^{\ell},\rho_{t}^{\ell})$ must be observed for any $\delta$. Consequently, any violation of this monotonically non-increasing behaviour of these two quantities is a signature of indivisibility as well as "backflow" of information (about the distinguishability between quantum states) from the environment back to the subsystem in question and thus, of non-Markovianity, and a degree of non-Markovianity may then be defined by the magnitude of such violations.
Model and methods— We demonstrate our results on the paradigmatic one-dimensional Ising spin chain in the presence of both transverse and longitudinal fields with open boundary conditions,
$\mathcal{H}_{I}=-J\sum_{j=1}^{N-1}\sigma^{z}_{j}\sigma_{j+1}^{z}-h_{x}\sum_{j=1}^{N}\sigma_{j}^{x}-h_{z}\sum_{j=1}^{N}\sigma_{j}^{z}\text{ .}$
When $h_{z}\!=\!0$, the model is integrable, with the Ising critical point at $J\!=\!h_{x}$ separating the symmetry-broken ferromagnetic/antiferromagnetic phases from the paramagnetic phase [30]. When $h_{z}\!\neq\!0$, the model is non-integrable and in the symmetry-broken phases (ferromagnetic or anti-ferromagnetic) exhibits confinement between kink-antikink excitations [31, 32, 33, 34] and consequently slow thermalization after a quench to this non-integrable regime [35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
For our simulations we use the matrix product state (MPS) formalism and second-order time-evolving block decimation (TEBD2) algorithm for simulating the real-time dynamics [45, 46, 47]. Our simulations were performed with the ITensors library in Julia [48, 49]. Most results shown here were obtained with Trotter time-steps $\tau\!=\!0.01$ seconds for our TEBD2 simulations (thus, our numerical errors are $\mathcal{O}(10^{-4})$), and we have checked that the results were unchanged with a smaller time-step of $\tau\!=\!0.002$ seconds. The total system size was fixed at $N\!=\!200$ sites, and results were independent of the other system sizes used for verification purposes. The MPS cutoffs were set at $10^{-9}$, and maximum bond-dimensions were set at $50$. The ground states were obtained with the density matrix renormalization group (DMRG) algorithm [46], also implemented with the same cutoffs as mentioned above.
In this work we have focused on investigating small subsystems comprising of a few spins embedded in the (translationally-invariant) bulk of a large spin chain. This is partly due to the fact that the computational costs of constructing reduced density matrices scale exponentially in the subsystem sizes, and in fact already for subsystems of only five spins this was evident to us in practice. This also partly served the purpose of going systematically in investigating the dynamics of subsystems starting up from the smallest possible one. We will see below that already for a four-spin subsystem, the non-Markovianity of its dynamics is quite insignificant, so our focus on small subsystems gets incentivized a posteriori. Perhaps most importantly, in contemporary ultracold atomic experiments, local site-resolved studies of quantum states has been made feasible and put to practice (see e.g. [50, 51]), which motivates putting the microscope on very small subsystems also theoretically.
Figure 1: Paramagnetic-to-ferromagnetic quench in mixed-field Ising spin chain: Trace distances (TD) as functions of time for various small subsystem RDMs with $\delta=1$ sec. (left), and $\delta=2$ sec. (right). The highly non-monotonic behaviour signifies strong non-Markovianity.
Figure 2: Degree of subsystem non-Markovianity for TD distances vs $\delta$ in the mixed-field Ising spin chain. Upper Row - for various small subsystems in paramagnetic to ferromagnetic quench (left), and ferromagnetic to paramagnetic quench (right). Lower Row - for 4-RDMs with varying longitudinal field (left) and transverse field (right).
Results— In Fig.1, we show results for the evolution of trace distance between reduced density matrices of subsystems composed of one to four spins (multiple spins in contiguous blocks), separated in time by $\delta\!=\!1$ and $2$ seconds for a quench starting from the paramagnetic ground state at $(J,h_{x},h_{z})\!=\!(0.2,1,0)$ to the ferromagnetic side $(J,h_{x},h_{z})\!=\!(1,0.1,0.5)$. Note that the latter parameters are deep in the non-integrable regime. The highly non-monotonic behaviour of the trace distances in Fig.1 is a telltale signature of strong non-Markovianity in the dynamics of these subsystems. Moreover, the oscillatory behaviour in trace distances appear to be rather persistent. This, we believe, is also a signature of very slow relaxation and anomalous thermalization, since it stands to reason that strong non-Markovianity should in general be an obstruction to fast and efficient relaxation. Note also that at any fixed time $t$, the trace distances between the temporally-separated states of the larger subsystems are strictly higher than those of the smaller subsystems : this is simply due to the contractivity of any distance measures under the partial trace operation.
A more revealing picture may be obtained by defining a simple degree of non-Markovianity as the cumulative magnitude of the revivals (increases) of the trace distances (other definitions of such a degree are possible of course). Fig.2 shows this degree, the upper row of which reveals two facts : I) no discernible mathematical pattern is seen for the case of paramagnetic-to-ferromagnetic quench except for the fact that this degree assumes noticeably smaller values for the larger subsystems of three and four spins, signifying that the dynamics of smaller subsystems of one and two spins are considerably more non-Markovian and further larger subsystems have progressively lessening non-Markovianity in their dynamics, and II) a negligible degree of non-Markovianity is seen in the case of ferromagnetic-to-paramagnetic quench (about $50$ times smaller than the opposite case), although a more systematic dependence on the subsystem sizes is seen. The subsystems’ dynamics in the latter case may thus be considered as effectively Markovian. We have seen similar results for paramagnetic-to-paramagnetic and ferromagnetic-to-ferromagnetic quenches as well. It is quite surprising that these latter quenches apparently induce effectively Markovian dynamics of subsystems (at least as per the behaviour of distance measures between quantum states), since it seems more natural that subsystems of closed quantum many-body systems should naturally exhibit some notable degree of non-Markovianity in far from equilibrium times for any parameter regimes because the process of equilibration of a closed quantum many body system is a complex process with various subsystems interacting with (and acting as "environments" for) each other in a common pursuit to attain a globally equilibrated state and the dynamics of a given subsystem (especially one that is sufficiently small compared to the total system size) plus its complement can not in general be expected to flow monotonously (i.e., a monotonous flow of information out of the subsystem to its environment) to equilibration in general, even more so because the environments themselves are also evolving dynamically and the interaction strengths between intra-subsystem spins, intra-environment spins and inter-subsystem-environment spins are all equal to each other. Yet, it is the direction of quenches and the consequent restructuring and mixing of the quantum states that may be the most decisive factor on whether the ensuing dynamics of subsystems is Markovian or otherwise.
The lower row of Fig.2 shows the dependence of this degree, for the case of a 4-spin subsystem (other subsystems show similar behaviour), on various values of longitudinal and transverse fields after quenching from the paramagnetic ground state. Here again, no discernible pattern is seen although a crude decrease of the degree is seen with increasing transverse fields.
Figure 3: First row - Non-monotonicities of TVD vs time for $\delta=1,2$ seconds for paramagnetic-to-ferromagnetic quench. Second Row - Degree of TVD non-monotonicities vs $\delta$ in paramagnetic-to-ferromagnetic quench (left), and ferromagnetic-to-paramagnetic quench (right). Third Row - Dependence of this degree on varying longitudinal field (left) and transverse field (right) for 4-RDMs in the former quench.
Remarkably, a more organized behaviour in the corresponding quantities is shown by the classical probability distributions furnished by the eigenvalues of the RDMs. Let us hypothesize for the moment that under CPTP maps acting on a pair of quantum states (density matrices), the distance between the corresponding probability vectors furnished by their respective eigenvalues also shows a monotonically non-increasing behaviour [[Althoughintuitivelyappealing, wearenotawareofanysuchresultsestablishedinquantuminformationtheorytothebestofourknowledge.Inclassicalinformationtheory, distancemeasuresbetweenprobabilitydistributionsalsoshowmonotonicallynon-increasingbehaviourunderclassicalprocessesandconsequentinequalitiesareknownasdata-processinginequalities, seee.g.][, whosequantumcounterpartsarethemonotonicinequalitiesgoverningdistancesbetweenquantumstatesunderquantumchannels/CPTPmaps.Butwhatclassical"dataprocessing"aspectsarecontainedwithinquantumchannelsactingonquantumstatesisasubtlequestioninquantuminformationtheory.]Cover1999]. If this hypothesis is true, could it be that any violation of this monotonicity under the action of a family of CPTP maps (dynamical maps) is also a signature of some aspect of non-Markovian dynamics of the underlying quantum states? [[Itisintriguingtopositthatthequantumdynamicsofsubsystemsalsocarry"classical"aspectsinsuchamannerthattheeigenvaluesofrespectivereduceddensitymatricesmightshowtheclassically-expectedmonotonicitiesfordistancemeasuresbetweenthemiftheunderlyingquantumdynamicsisMarkovian, andaviolationofthosemonotonicitieswhentheunderlyingquantumdynamicsisnon-Markovianasperquantummeasures.Itistemptingtoregardthisasaclassicalcharacteristicofquantumnon-Markovianitybutinasensedifferentfrom][.Itwouldbeworthinvestigatingthissystematicallyinfuture.]Milz2020]. Be that as it may, if the above hypothesis is true, then a degree of violation of said monotonicity may be defined again by the cumulative magnitude of the revivals (increases) of the employed distance measure. Fig.3 shows the non-monotonicities of the total variation distances defined previously and the corresponding "degree of non-monotonicities" for the paramagnetic-to-ferromagnetic quench, which shows several remarkable facts. Firstly, in the second row, the degree for any considered subsystem is strongly oscillatory but decreasing on an average with increasing temporal separation $\delta$. Secondly, the degrees for ferromagnetic-to-paramagnetic quench (right, second row) are about $\sim 30$ times smaller than those for the paramagnetic-to-ferromagnetic quench (left, second row). Thirdly, for paramagnetic-to-ferromagnetic quench (left, second row), the $\delta$-values corresponding to the visible maxima and minima are nearly identical for all considered subsystems (notice however that the maxima corresponding to 1-RDMs are slightly shifted to the left relative to those corresponding to the other subsystems), and they appear to nearly coincide with each other already for subsystems as small as three and four spins. Note however that this degree at any given $\delta$ is larger in value for larger subsystems, in contrast with the degrees of non-Markovianity shown in Fig.2. Clearly then, this degree can not be a measure of non-Markovianity of quantum states in the usual sense. Fourthly, in the third row of Fig.3 a general increase of these degrees with increasing $h_{z}$ is seen (for most values of $\delta$), whereas a rather drastic reduction in their oscillatory amplitude is seen with increasing $h_{x}$. Once again, quite remarkably the maxima and minima of these degrees at varying values of $h_{z}$ and $h_{x}$ occur at nearby $\delta$-values, with relative shifts becoming more prominent at higher $\delta$-values (especially for the integrable case $h_{z}=0$). Moreover, there seem to exist remarkable timescales associated with the behaviour of TVD between the eigenvalues of considered RDMs : a timescale of $\sim 0.78(\pm 0.02)$ seconds in evolution times $t$ between the minima in the first row of Fig.3, which is largely independent of the subsystem size and of the temporal-separation $\delta$ (we checked this for several $\delta$ values), and a second timescale of $\sim\{1.55,1.6\}$ seconds in the temporal-separation $\delta$ between the maxima (or between the minima) emerges in the degrees of TVD non-monotonicities (second row) which also is largely independent of the subsystem size (additionally, even with varying transverse and longitudinal fields there is a timescale of approximately the same value but in a cruder fashion).
[…]
We remarked earlier that the magnetically-ordered regime with non-zero longitudinal field is known to exhibit slow thermalization dynamics. This is mainly due to the existence of confining potential between certain excitations that separate domain walls [31, 32, 33, 34], as well as due to the presence of scars (non-thermalizing eigenstates) deep in the spectrum whose presence is related to and indeed enhanced by the presence of confinement [59, 41]. The strength of the confining potential is directly proportional to the longitudinal field and we have seen in this article that, while the trace distance degrees signifying subsystem non-Markovianity did not show sufficiently differential behaviour with increasing values of the longitudinal field, the degrees of non-monotonicities of total variation distance between eigenvalues of subsystem RDMs showed marked increase with increasingly higher longitudinal fields. This intriguing connection would be worth understanding more precisely in future. Additionally, the asymmetry between the two directions of quenches, seen in this work in the context of subsystem non-Markovianity, was also noted previously in the context of dynamical quantum phase transitions and differentiating dynamical behaviour was seen in quantum mutual information (between small subsystems) and in leading members of the entanglement spectrum [60].
Concluding remarks— By numerically investigating the dynamical behaviour of trace distances between temporally-separated reduced density matrices corresponding to small subsystems of a closed quantum spin system in one dimension, and with a viewpoint of looking at these subsystems as open quantum systems, we have presented numerical evidence of non-Markovianity of these subsystems’ dynamics and the dependence of this characteristic on various factors at play. Notably, it was demonstrated that the direction (in parameter space) of quenching is a major factor for noticeable non-Markovianity, and that smaller subsystems showcase larger non-Markovianity in their dynamics. We have also revealed a remarkably systematic dynamical behaviour in the measure of distance between the corresponding eigenvalues of the considered reduced density matrices, the interpretation and explanation of which we can not offer at this time and will be left for future work. We hope this sort of a viewpoint can be put to practice more widely by applying some of the central tools and concepts of open quantum systems theory [22, 23] onto uncovering and understanding many facets of non-equilibrium dynamics of isolated quantum many-body systems. While we focused in this article on the Ising spin model for simplicity and the paradigmatic value held by this model, we have observed signatures of subsystem non-Markovianity in quench dynamics of Heisenberg spin systems also, albeit they were most pronounced in quenches from paramagnetic to magnetically ordered phases in those systems as well, hence our focus on the simpler Ising spin chain in this article. In a forthcoming article we will present results on the dynamical behaviour of other quantum informational quantities that accompany subsystem non-Markovianity, as well as an example of another simple spin system where, unlike the Ising model in this work, a quench from a magnetically ordered to a paramagnetic side (where the latter is topological in nature) also shows pronounced non-Markovianity. Further important questions that need to be addressed include the effects of performing measurements on subsystem non-Markovianity, whether effective spin-boson type models or collision model descriptions can be constructed to mimic the open system dynamics of subsystems of closed many-body systems, a characterization of the properties of the dynamical maps themselves, and whether other frameworks for characterizing and simulating non-Markovianity can provide additional or deeper insights. Work is in progress in some of these directions and we hope to be able to report on them in future.
Acknowledgements— Financial support from the Department of Atomic Energy, India is gratefully acknowledged.